A student found an unknown white powder in his refrigerator. His mom tells him that it is either baking soda or baking powder. He decided to try to figure out what it is through analytical chemistry.


1. Baking soda is simply sodium bicarbonate (NaHCO3), so to begin the student set out to compare the two by percent by mass. What is the percent Na, H, C, and O in sodium bicarbonate? Show your work.

a.    %Na


b.    %H


c.    %c


d.    %O



2.The student then looked up the difference between the two on the internet. It seems baking powder is a mixture of multiple substances (NaHCO3, KHCO3, KC4H5O6, etc), while baking soda is a pure compound. Based on this would baking powder have a higher, lower, or similar percent sodium by mass? Explain in simple terms.



3.The student predicts that the white powder IS baking soda (NaHCO3 – a pure substance). Do you agree with the student, or disagree and why?


1
Expert's answer
2021-02-24T04:15:46-0500

1. a. %Na = 23 / (23+1+12+16×3) = 27,38%

b. %H = 1/(23+1+12+16×3) = 1,19%

c. %C = 12/(23+1+12+16×3) = 14,29%

d. %O = 16×3/(23+1+12+16×3) = 57,14%


2. Baking powder has a lower percent sodium by mass, since baking powder has some other compounds in its composition, so to count sodium percentage we have to divide it's mass by the whole mixture mass, not only by sodium bicarbonate mass. Roughly speaking other ingredients in baking powder "take" some percents of mass, that's why percentage of sodium in it is lower, than in pure baking soda.


3. I disagree, since the majority of compounds of sodium and potassium is white, baking soda and baking powder are white, in particular